# ChatStitch: Visualizing Through Structures via Surround-View Unsupervised Deep Image Stitching with Collaborative LLM-Agents

### v1

@@ Abstract
Collaborative perception has garnered significant attention for its ability to enhance the perception capabilities of individual vehicles through the exchange of information with surrounding vehicle-agents. However, existing collaborative perception systems are limited by inefficiencies in user interaction and the challenge of multi-camera photorealistic visualization. To address these challenges, this paper introduces ChatStitch, the first collaborative perception system capable of unveiling obscured blind spot information through natural language commands integrated with external digital assets. To adeptly handle complex or abstract commands, ChatStitch employs a multi-agent collaborative framework based on Large Language Models. For achieving the most intuitive perception for humans, ChatStitch proposes SV-UDIS, the first surround-view unsupervised deep image stitching method under the non-global-overlapping condition. We conducted extensive experiments on the UDIS-D, MCOV-SLAM open datasets, and our real-world dataset. Specifically, our SV-UDIS method achieves state-of-the-art performance on the UDIS-D111we extract a subset of data from the original UDIS-D dataset for the multi-image stitching experiment. dataset for 3, 4, and 5 image stitching tasks, with PSNR improvements of 9%, 17%, and 21%, and SSIM improvements of 8%, 18%, and 26%, respectively.
@@ I Introduction
Perception plays a crucial role in intelligent unmanned systems, where its primary goal is to extract pertinent information from the surrounding environment and convey it to subsequent tasks such as mapping, planning, and control. With the continuous advancements in deep learning, individual perception has made significant strides across various perception tasks, including image stitching[1, 2, 3, 4], object detection[5, 6, 7, 8], and semantic segmentation[9, 10, 11, 12]. Despite the immense potential exhibited by individual perception, this approach often encounters issues such as occlusions, primarily due to the limited viewpoint of a single agent. To address this challenge, collaborative perception[13] has garnered widespread attention from both the academic and industrial communities in recent years. The objective of collaborative perception is to enhance the perception capabilities of individual vehicles by exchanging complementary information with surrounding agents.
@@
Recent efforts have made valuable contributions both in the development of high-quality real or simulated datasets and in devising effective solutions to realize collaborative perception[14, 15, 16, 17, 18]. However, existing collaborative perception approaches predominantly focus on the exchange of information between agents, with a notable absence of human involvement. On one hand, current collaborative schemes lack responsiveness to human directives[19, 20, 21]; on the other hand, these approaches primarily concentrate on enhancing the effectiveness of perception tasks[22, 23, 24], such as object detection, and often overlook the direct intuitive experiences of humans in photo-realistic environments.
To address these needs, we introduce ChatStitch, the first collaborative perception system capable of revealing obscured blind spot information through natural language commands integrated with external digital assets. To utilize ChatStitch, users simply engage in dialogue with the system, issuing commands in natural language without the need to partake in the intermediate perception steps. This process is illustrated in Fig. 1.
To adeptly handle complex or abstract commands, ChatStitch employs a multi-agent collaborative framework based on Large Language Models (LLMs). The core concept revolves around utilizing multiple LLM-agents, each designated with a specific role, to decompose the overall collaborative perception requirements into distinct tasks. This reflects the functional segmentation typically established within technological frameworks. This workflow offers two primary advantages. Firstly, the Large Language Models are capable of processing human natural language commands, allowing for intuitive and dynamic adjustments to the collaborative perception displays, thereby achieving precise tuning and feedback. Secondly, the collaborative framework enhances the effectiveness and accuracy of collaborative perception by allocating specific tasks among specialized agents, thus expanding the boundaries of collaborative perception capabilities.
@@
To achieve the most intuitive perception for humans, we propose SV-UDIS in ChatStitch, which stitches multiple surround view images in driving scenarios into a panoramic image. The current state-of-the-art (SoTA) unsupervised image stitching algorithms[3, 25] mainly focus on two-image stitching tasks. Although they can also perform multi-image stitching, they are not suitable for surround view images in driving scenarios. On one hand, in these scenarios, the overlapping areas between the images to be stitched are extremely small, resulting in severe stretching at the edges of the warped images. On the other hand, only adjacent images have overlapping areas, so there is no common reference image that has shared view areas with all other images, as shown in Fig. 2(b). For clarity, we define the above surround-view stitching task as a multi-image stitching problem under the non-global-overlapping condition. A detailed definition of this problem can be found in Section IV-A; more discussion on multi-image stitching please refer to our supplementary materials. Our proposed SV-UDIS method, through masked cylindrical projection, rectangular warping constraints, and motion propagation, achieves multi-image stitching under the non-global overlapping condition, providing humans with intuitive visual effects.
@@
  • We develop an innovative framework, named ChatStitch, which is the first collaborative perception system capable of unveiling obscured blind spot information through natural language commands integrated with external digital assets.
  • We propose SV-UDIS, the first surround-view unsupervised deep image stitching framework under the non-global-overlapping condition, which stitches common surround images in driving scenarios into a panorama, providing intuitive visual effects.
@@
Method # Photo- realistic Dim. Multi- camera Editable External assets Language Open- source Stitching
@@
II Related work
Unsupervised image stitching. Image stitching technology[48] combines multiple images with overlapping areas to generate an image with a larger field of view. In recent years, learning-based image stitching methods[49] have emerged. Among them, due to the difficulty in obtaining real stitched labels, unsupervised stitching methods[2, 3, 50, 51] represented by UDIS[2] and UDIS2[3] are more popular than supervised methods[52, 53, 54, 55]. However, few have considered multi-image stitching under the non-global-overlapping condition, which is common in driving scenarios, such as the camera configurations used in the NuScenes[56], Waymo[57] and PandaSet[58] datasets, where all images are horizontally arranged with extremely low overlapping areas. In this paper, we propose SV-UDIS, a multi-image stitching algorithm for non-global-overlapping situations.
Large Language Model and multi-agent framework. Large Language Models are designed to integrate vast amounts of data for understanding, generating, and responding to human language. Since Google introduced the Transformer model [59], a plethora of models and methods based on it have emerged, particularly exemplified by the Large Language Models such as GPT[60] and BERT[61]. Building on LLMs, many efforts have begun utilizing rich multi-agent collaborative pattern components[24], enabling agents to function optimally in their respective roles across various domains. In our work, we explore the use of a multi-agent framework to accomplish collaborative perception tasks and to display the intuitive results of such collaboration.
III Multi-Agent Collaborative Framework
ChatStitch analyzes complex and abstract commands in human-machine collaboration, providing users with photorealistic collaborative perception visualization videos. Please refer to Fig. 3. The challenge of directly applying a single LLM-agent lies in its struggle with multistep reasoning and cross-referencing [24]. Drawing from this insight, we employ a multi-agent collaborative framework with LLM agents, where each agent is responsible for managing a specific module of the collaborative perception task.
@@ III-A Agents’ Functionality
Foreground Rendering. The Foreground Rendering Agent combines viewpoint data and 3D asset information to render partners in a collaborative perception scenario. Its LLM module receives vehicle IDs and activates a specific function, which uses these parameters to call the instant-ngp interface[62] for rendering partners’ images.
@@ III-B Workflow
ChatStich is composed of multiple agents that collaboratively produce panoramic images of a scene based on human instructions. The Task Management Agent breaks down complex human instructions into simpler task directives and delivers them to subsequent other agents. These agents work in two separate processes: Background Stitching and Foreground Rendering.
@@
The final images from Foreground Rendering and Background Stitching are integrated to achieve a photo-realistic result for collaborative perception. Thanks to the Task Management Agent’s ability to store instructions and parameters for other agents, it enables scene modifications through multiple rounds of instructions, enhancing dynamic interaction and responsiveness in the scene.
IV SV-UDIS
In this section, we propose SV-UDIS, the first surround-view unsupervised deep image stitching method under the non-global-overlapping condition. we first describe the definition of the non-global-overlapping condition, and then introduce the overall framework of SV-UDIS, as shown in Fig. 4.
@@ IV-A Problem Definition
In this paper, we aim to stitch $N$ adjacent images captured by the surround view cameras into a panoramic image, with the camera configuration identical to that of MCOV-SLAM[26]. For a set of surround view image data $\{I_{i}\}_{i=1,...,N},N>3$ with the resolution of $W\times H$ collected at a certain moment, we define the non-global-overlapping condition (shown in Fig. 2(b)) as: for any image $I_{i}$ captured by camera $i$, there is and only exists overlapping areas with images captured by adjacent cameras. Represented as:
@@ IV-B Unsupervised Multi-Image Stitching
The proposed SV-UDIS framework, as shown in Fig. 4, mainly includes three stages: Masked cylindrical projection and feature extraction, multi-image warping, and multi-image composition. In the first stage, we perform cylindrical projection on the input $N$ images, generate projected images and projected masks, and extract multi-scale features. In the second stage, we take two adjacent images as a group, regress their features, and generate the mesh warping of the target image for each group. Then, using our proposed motion propagation strategy, we calculate the mesh motion between the edge image and the central image, thereby generating aligned images and masks through TPS transformation[3]. In the third stage, we predict the image fusion masks through a UNet-like network, thereby generating the final stitched image. A more detailed processing procedure are reported in the supplementary materials.
Masked cylindrical projection. In this paper, we aim to stitch multiple images with a large horizontal field of view. To maintain the spatial constraints and visual consistency of the images, we need to process the images with cylindrical projection, i.e., project the images from each camera onto the cylindrical surface, and then unfold the cylindrical surface to achieve the projection deformation of each image. The impact of whether to perform cylindrical projection on the multi-image stitching task with a large horizontal field of view is shown in Fig. 2. A detailed procedure of cylindrical projection can be found in our supplementary material.
@@
Multi-agent collaboration language command category
Show Unshow Select Abstract
0.704 0.211 0.368 0.462
0.963 0.947 0.789 0.923
[…]
@@
TABLE III: Quantitative comparison of multi-image stitching on UDIS-D[2] and MCOV-SLAM[26] datasets. Red indicates the best results. The symbol denotes that we extract a subset of data from the original UDIS-D[2] dataset for the multi-image stitching experiment.
To mitigate the issue of shape stretching at the edges and to ensure a more natural visual effect in the stitched image, we propose the rectangular warping constraints. The basic idea behind these constraints is that the warped target image should be as close to a rectangle as possible, avoiding severe shape stretching and scale scaling. We design a loss function to implement these constraints, as shown in Eq. 4 , which is composed of three parts: rectangle shape constraint, rectangle size constraint, and grid folding penalty. For the specific meanings and details of each term in the formula, please refer to our supplementary materials.

### v2

@@ Abstract
Surround-view perception has garnered significant attention for its ability to enhance the perception capabilities of autonomous driving vehicles through the exchange of information with surrounding cameras. However, existing surround-view perception systems are limited by inefficiencies in unidirectional interaction pattern with human and distortions in overlapping regions exponentially propagating into non-overlapping areas. To address these challenges, this paper introduces ChatStitch, a surround-view human-machine co-perception system capable of unveiling obscured blind spot information through natural language commands integrated with external digital assets. To dismantle the unidirectional interaction bottleneck, ChatStitch implements a cognitively grounded closed-loop interaction multi-agent framework based on Large Language Models. To suppress distortion propagation across overlapping boundaries, ChatStitch proposes SV-UDIS, a surround-view unsupervised deep image stitching method under the non-global-overlapping condition. We conducted extensive experiments on the UDIS-D, MCOV-SLAM open datasets, and our real-world dataset. Specifically, our SV-UDIS method achieves state-of-the-art performance on the UDIS-D111we extract a subset of data from the original UDIS-D dataset for the multi-image stitching experiment. dataset for 3, 4, and 5 image stitching tasks, with PSNR improvements of 9%, 17%, and 21%, and SSIM improvements of 8%, 18%, and 26%, respectively.The code is available at https://github.com/lhlawrence/ChatStitch.
Index Terms:
surround view, image stitching, human-machine interaction.
@@ I Introduction
Surround-view Perception plays a crucial role in autonomous driving, where its primary goal is to extract pertinent information from the surrounding environment and convey it to subsequent tasks such as mapping, planning, and control[1]. With the continuous advancements in large language models(LLMs), the fusion of human-centric design and artificial intelligence capabilities has opened up new possibilities for next-generation autonomous vehicles that go beyond traditional transportation[2][3].
@@
However, existing visionLLM systems typically rely on manually preset parameters or unidirectional command inputs[4][5][6]. Interaction is limited to the initial command input, and the correction process is fully automated without iterative optimization, resulting in low interaction efficiency. To enhance system efficiency, bidirectional interaction, including real-time user feedback on correction outcomes, is necessary. Additionally, in the context of surround-view perception, the classic task of panoramic image stitching is challenged by the constraints of vehicle layout. The overlapping regions between different images are often too small, which causes distortion propagation across these overlapping boundaries. The current state-of-the-art (SoTA) unsupervised image stitching algorithms[7, 8] mainly focus on two-image stitching tasks. Although they can also perform multi-image stitching, they are not suitable for surround view images in driving scenarios. On one hand, in these scenarios, the overlapping areas between the images to be stitched are extremely small, resulting in severe stretching at the edges of the warped images. On the other hand, only adjacent images have overlapping areas, so there is no common reference image that has shared view areas with all other images, as shown in fig. 2(b). For clarity, we define the above surround-view stitching task as a multi-image stitching problem under the non-global-overlapping condition. A detailed definition of this problem can be found in section IV-A; more discussion on multi-image stitching please refer to our supplementary materials.
@@
To address these needs, we introduce ChatStitch, a surround-view human-machine co-perception system capable of unveiling obscured blind spot information through natural language commands integrated with external digital assets. To utilize ChatStitch, users simply engage in dialogue with the system, issuing commands in natural language without the need to partake in the intermediate perception steps. This process is illustrated in fig. 1.
To dismantle the unidirectional interaction bottleneck, ChatStitch implements a cognitively grounded closed-loop interaction multi-agent framework based on LLMs. The core concept revolves around utilizing multiple LLM-agents, each designated with a specific role, to decompose the overall surround-view human-machine co-perception perception requirements into distinct tasks. This reflects the functional segmentation typically established within technological frameworks. This workflow offers two primary advantages. Firstly, the Large Language Models are capable of processing human natural language commands, allowing for intuitive and dynamic adjustments to the perception displays, thereby achieving precise tuning and feedback. Secondly, the closed-loop framework enhances the effectiveness and accuracy of perception by allocating specific tasks among specialized agents, thus expanding the boundaries of surround-view perception capabilities.
To suppress distortion propagation across overlapping boundaries, We propose SV-UDIS within ChatStitch, a technical module corresponding to an agent in the ChatStitch framework, designed to stitch multiple surround-view images captured in driving scenarios into a seamless panoramic image. Our proposed SV-UDIS method, through masked cylindrical projection, rectangular warping constraints, and motion propagation, achieves multi-image stitching under the non-global overlapping condition, providing humans with intuitive visual effects.
@@
  • We develop a cognitively grounded closed-loop interaction multi-agent framework, named ChatStitch, which is a surround-view human-machine co-perception system capable of unveiling obscured blind spot information through natural language commands integrated with external digital assets.
  • We propose SV-UDIS within ChatStitch, a surround-view unsupervised deep image stitching method under the non-global-overlapping condition, which stitches common surround images in driving scenarios into a panorama, providing intuitive visual effects.
@@
II Related work
Unsupervised image stitching. Image stitching technology[11] combines multiple images with overlapping areas to generate an image with a larger field of view. In recent years, learning-based image stitching methods[12] have emerged. Compared to traditional methods[13, 14, 15, 16] based on sparse geometric features[17, 18, 19], learning-based stitching methods have shown better robustness and faster stitching speed, as they eliminate the endless design of geometric features. Among them, due to the difficulty in obtaining real stitched labels, unsupervised stitching methods[9, 7, 20, 21] represented by UDIS[9] and UDIS2[7] are more popular than supervised methods[22, 23, 24, 25]. However, few have considered multi-image stitching under the non-global-overlapping condition, which is common in driving scenarios, such as the camera configurations used in the NuScenes[26], Waymo[27] and PandaSet[28] datasets, where all images are horizontally arranged with extremely low overlapping areas. In this paper, we propose SV-UDIS, a multi-image stitching algorithm for non-global-overlapping situations.
Large Language Model and multi-agent framework. Large Language Models are designed to integrate vast amounts of data for understanding, generating, and responding to human language. Since Google introduced the Transformer model [29], a plethora of models and methods based on it have emerged, particularly exemplified by the Large Language Models such as GPT[30] and BERT[31]. Building on LLMs, many efforts have begun utilizing rich multi-agent collaborative pattern components[32], enabling agents to function optimally in their respective roles across various domains. In our work, we explore the use of a multi-agent framework to accomplish surround-view perception tasks.
Method Photo- realistic Dim. Multi- camera Editable External assets Language Open- source Stitching
@@ II Related work
III Multi-Agent Framework
ChatStitch analyzes complex and abstract commands in human-machine co-perception system, providing users with surround-view perception visualization videos. Please refer to fig. 3. Given the complexity of human language and the abstract nature of human-machine co-perception tasks, our system is designed to flexibly handle specific aspects or multiple components of these tasks. The challenge of directly applying a single LLM-agent lies in its struggle with multi-step reasoning and cross-referencing [32]. Drawing from this insight, we employ a multi-agent framework with LLM agents, where each agent is responsible for managing a specific module of the perception task.
@@ III-A Agents’ Functionality
Foreground Rendering. The Foreground Rendering Agent combines viewpoint data and 3D asset information to render partners in a surround-view perception scenario. Its LLM module receives vehicle IDs and activates a specific function, which uses these parameters to call the instant-ngp interface[60] for rendering partners’ images.
@@ III-B Workflow
ChatStich is composed of multiple agents that collaboratively produce surround-view images of a scene based on human instructions. The Task Management Agent breaks down complex human instructions into simpler task directives and delivers them to subsequent other agents. These agents work in two separate processes: Background Stitching and Foreground Rendering.
@@
The final images from Foreground Rendering and Background Stitching are integrated to achieve a photo-realistic result for surround-view perception. Thanks to the Task Management Agent’s ability to store instructions and parameters for other agents, it enables scene modifications through multiple rounds of instructions, enhancing dynamic interaction and responsiveness in the scene. The closed-loop framework enhances the effectiveness and accuracy of perception by allocating specific tasks among specialized agents, thus expanding the boundaries of surround-view perception capabilities.
IV SV-UDIS within Background Stitching Agent
In this section, We propose SV-UDIS within ChatStitch, a technical module corresponding to an agent in the ChatStitch framework, designed to stitch multiple images captured in driving scenarios into a seamless panoramic image under the non-global-overlapping condition. we first describe the definition of the non-global-overlapping condition, and then introduce the overall framework of SV-UDIS, as shown in fig. 4.
@@ IV-A Problem Definition
In this paper, we aim to stitch $N$ adjacent images captured by the surround-view cameras into a panoramic image, with the camera configuration identical to that of MCOV-SLAM[10]. For a set of surround view image data $\{I_{i}\}_{i=1,...,N},N>3$ with the resolution of $W\times H$ collected at a certain moment, we define the non-global-overlapping condition (shown in fig. 2(b)) as: for any image $I_{i}$ captured by camera $i$, there is and only exists overlapping areas with images captured by adjacent cameras. Represented as:
@@
Our goal is to construct a self-supervised stitching method $\mathbf{S}(\cdot)$ with learnable parameters $\theta$, which takes $N$ images satisfying the non-global-overlapping condition described in eq. 1 as input, and stitches them to generate a panoramic image $I_{sv}$. It is defined as
$\mathbf{S}(I_{1},...,I_{N};\theta)\to I_{sv}$ (3)
@@ IV-B Unsupervised Multi-Image Stitching
The proposed SV-UDIS framework, as shown in fig. 4, mainly includes three stages: Masked cylindrical projection and feature extraction, multi-image warping, and multi-image composition. In the first stage, we perform cylindrical projection on the input $N$ images, generate projected images and projected masks, and extract multi-scale features. In the second stage, we take two adjacent images as a group, regress their features, and generate the mesh warping of the target image for each group. Then, using our proposed motion propagation strategy, we calculate the mesh motion between the edge image and the central image, thereby generating aligned images and masks through TPS transformation[7]. In the third stage, we predict the image fusion masks through a UNet-like network, thereby generating the final stitched image.
Masked cylindrical projection and feature extraction. In this paper, we aim to stitch multiple images with a large horizontal field of view. To maintain the spatial constraints and visual consistency of the images, we need to process the images with cylindrical projection, i.e., project the images from each camera onto the cylindrical surface, and then unfold the cylindrical surface to achieve the projection deformation of each image. The impact of whether to perform cylindrical projection on the multi-image stitching task with a large horizontal field of view is shown in fig. 2.
$P=K^{-1}p$ (4)
$\lambda=u,\varphi=\arctan v$ (5)
$\Rightarrow$ $X=\sin\lambda,Y=v,Z=\cos\lambda$
$\Rightarrow$ $p=KP$
@@
To mitigate the issue of shape stretching at the edges and to ensure a more natural visual effect in the stitched image, we propose the rectangular warping constraints. The basic idea behind these constraints is that the warped target image should be as close to a rectangle as possible, avoiding severe shape stretching and scale scaling. We design a loss function to implement these constraints, as shown in eq. 7 , which is composed of three parts:rectangle shape constraint $\ell_{shape}$, rectangle size constraint $\ell_{size}$, and grid folding penalty $\ell_{fold}$.
For the grid folding penalty, the warped grid should not exhibit folding phenomena, that is, the relative positional relationship among all grid edges — up, down, left, and right — should remain unchanged. Therefore, we control the orientation of the grid edge vector $\vec{e}_{hor}$ and $\vec{e}_{ver}$ through the $RELU$ function $\sigma(\cdot)$. It is specifically defined as follows:
Multi-Images Composition. We adopt the unsupervised seamless composition module proposed by UDIS2[7] for image composition, and extend it to multi-image stitching. For the warped images and masks, we also take two adjacent images as a pair, and then predict the composition masks for each pair of images. At this point, except for the two images located at the edge of the field of view, each other image has two composition masks. Multiplying them together gives us the final composition masks shown in fig. 4. Finally, the final stitched image can be generated using the warped images and the final composition masks.
Figure 6: Result under an abstract command.
V Experiments
V-A Experimental settings
ChatStitch. For system-level perception tasks involving Large Language Models, we conducted tests on our own real-world datasets. We aligned the positioning data from integrated navigation with panoramic image data based on GPS time, selecting a 15-second segment of typical vehicle occlusion and revelation for testing. All our Large Language Models utilized GPT-4o.
TABLE II: Quantitative comparison of multi-image stitching on UDIS-D[9] and MCOV-SLAM[10] datasets. Red indicates the best results. The symbol denotes that we extract a subset of data from the original UDIS-D[9] dataset for the multi-image stitching experiment.
@@ V-A Experimental settings
V-B Results
V-B1 System Results
Complex command. We provided the system with detailed information about the type, size, and location of vehicles. fig. 5 displays the final results. From this, we can observe that: 1) all the information requested in the complex commands has been successfully addressed, and 2) the intuitive effects of potential collaborative targets are visible even behind obstructing buildings.
Comparison with Visual Programming(VisProg)[64]. We utilized VisProg object replacement operation to remove occlusions and compared the results with our method, as shown in fig. 7. ChatStitch accurately performed perspective adjustment and localization of occluded objects, whereas VisProg consistently produced incorrect results. The reason for this discrepancy is that VisProg does not inherently support perspective, and is unable to obtain information about occluded objects.ct instructions.
Figure 7: Qualitative comparison with Visual Programming.
V-B2 Stitching Results
Quantitative comparison. We first conduct multi-image stitching comparison with UDIS2[7] on the two datasets, calculating PSNR and SSIM metrics to compare the image warping and alignment performance. The results are shown in table II(b), where $I_{3\times 3}$ represents the identity matrix, indicating no warping or alignment of the target image, serving as a reference. The evaluation results are divided into three levels based on performance, as described in UDIS[9] and UDIS2[7]. Next, we conduct two-image stitching comparison with other classic methods[14, 65, 15, 9] on these two datasets. The results are shown in table III.
TABLE III: Quantitative comparison of two-image stitching on the UDIS-D[9] and MCOV-SLAM[10] datasets. E, M, H, and A denote Easy, Moderate, Hard, and Average, respectively. Red indicates the best results.